Show that the relation p = {(a,b) | a -b is an integer} on the set of real numbers R is equivalence relation


1
Expert's answer
2021-08-16T13:51:30-0400

(I). Since "a-a=0" and is an even integer. "\left(a,\:a\right)\in R" . Therefore, R is reflexive.

(II). If "\left(a-b\right)" is even then "\left(b-a\right)" is also even. Hence, "\left(a-b\right)\in R" and "\left(b,\:a\right)\in R" . The relation indicate R is symmetric.

(III). If "\left(a,\:b\right)\in R" , "\left(b,\:c\right)\in R" , then "\left(a-b\right)" is even, "\left(b-c\right)" is even. Hence, "\left(a-b+b-c\right)=\left(a-c\right)" and "\left(a,\:c\right)\in R" . This shows that R is transitive.


Since R is reflexive, symmetric and transitive, it is an equivalence relation.